In a test of a​ gender-selection technique, results consisted of 264 baby girls and 19 baby boys. Based on this​ result, what is the probability of a girl born to a couple using this​ technique? Does it appear that the technique is effective in increasing the likelihood that a baby will be a​ girl?

The probability that a girl will be born using this technique is approximately _______ .

​(Type an integer or decimal rounded to three decimal places as​ needed.)

Does the technique appear effective in improving the likelihood of having a girl​ baby? yes or no
Answer #1

Total number of children = 264+19 =283

Number of baby girls = 264

Number of baby boys = 19

The probability that a girl will be born using this techinque = 264/283 = 0.9329 = 0.933 (rounding to three decimal places)

The probability that a baby girl is born is 0.933 which has improved from 0.5.

YES the technique appear to be effective in improving likelihood of having a girl baby.